# Single-pixel edge enhancement of object via convolutional filtering with localized vortex phase

### v1

@@ 3.2 Fidelity estimation of super-pixel method to generate localized vortex phase
The modulation accuracy of the super-pixel method between the target field $E_{\text{out}}$ and the desired field $E_{\text{desire}}$ can be quantified by calculating their fidelity $F$ as shown below:
$F=\frac{|E_{\text{desire}}^{*}\cdot E_{\text{out}}|^{2}}{|E_{\text{desire}}|^{2}|E_{\text{out}}|^{2}},$ (4)
where the two-dimensional distributions $E_{\text{out}}$ and $E_{\text{desire}}$ are reshaped into one column and their inner product is calculated in the numerator, and *** denotes the complex conjugate. $F=1$ indicates that the target fields and the desired fields are the same, whereas $F=0$ means they are completely different.
We conducted a numerical simulation to estimate the fidelity between the target field generated by the super-pixel method and the desired localized vortex phase with the amplitude functions, $(c/r)^{2}$, $c/r$, $\sqrt{c/r}$ and $ce^{-r^{2}/w^{2}}$, where $c$ denotes amplitude coefficient. The optimizing parameters to obtain the best fidelity in the super-pixel method are the amplitude coefficient $c$ and the aperture radius $R$. For the Gaussian amplitude function, an additional parameter $w$ needs to be optimized and here we used $w=100$ µm, which gives the highest fidelity of 0.89. The simulation conditions remain the same as those in Section 3.1, with an additional parameter, a focal length of the lens setting at 300 mm.
Figures 7 (a-d) show the simulation results of the fidelity between the ideal and the generated localized vortex phases with different $a(\bm{r})$: $(c/r)^{2}$, $c/r$, $\sqrt{c/r}$ and $ce^{-r^{2}/w^{2}}$, respectively. The horizontal and vertical axes in Fig. 7 show the aperture radius $R$ normalized by aperture position $\rho$ and the amplitude coefficient $c$ of $h(\bm{r})$, respectively. Table 1 shows summary of the maximum fidelities and the corresponding optimized parameters. The Gaussian amplitude function $e^{-r^{2}/w^{2}}$ provides fidelity $0.89$, whereas $\frac{1}{r^{2}}$, $\frac{1}{r}$ and $\frac{1}{\sqrt{r}}$ yield fidelity values below $0.6$ (refer to Table 1), because these power functions increase rapidly as they approach to the origin, resulting in low fidelity with the finite amplitudes in the super-pixel method. Even if the fidelity to the desired function is low, however, edge enhancement might be possible if a certain degree of localization is achieved by the amplitude function. Thus, in the next section, we use the amplitude functions $\frac{1}{r^{2}}$ and $e^{-r^{2}/w^{2}}$ which successfully performed edge enhancement in Fig. 6, to evaluate our single-pixel method via convolutional filtering.
Figure 7: (a-d) The fidelity of localized vortex phase with amplitude functions:$\frac{1}{r^{2}}$, $\frac{1}{r}$, $\frac{1}{\sqrt{r}}$ and $e^{-r^{2}/w^{2}}$ ($w=100$ µm) for different amplitude coefficients $c$ and aperture radius, $R$ normalized by aperture position $\rho$.
@@
Parameter $a(\bm{r})$ $\frac{1}{r^{2}}$ $\frac{1}{r}$ $\frac{1}{\sqrt{r}}$ $e^{-r^{2}/w^{2}}$ Max. fidelity, $F$ 0.02 0.45 0.57 0.89 Radius, $R/\rho$ 2.5 2.6 2 0.7 Amp. coefficient, $c$ 42 µm 100 µm 100 µm 9
@@
Now, we evaluate edge enhancement capability of our proposed single-pixel imaging method with the localized amplitude functions, $\frac{1}{r^{2}}$ and $e^{-r^{2}/w^{2}}$. We first prepared the binary hologram on the DMD to generate the shifted localized vortex phase $h(\bm{r}^{\prime}-\bm{r})$ by using the super-pixel method with $n=4$. By following the principles of Section 2.1, we calculated Fourier transform of the object light multiplied with the hologram on the DMD and then obtained optical power at the center point of the aperture, $(x,y)=(-\rho,n\rho)$, to reconstruct the two-dimensional edge enhanced image. The simulation conditions remain the same as those in Section 3.1.
Figure 8: (a) Distribution of PAO, (b) and (c) Edge enhancement via proposed method using optical vortex with the amplitude function $1/r^{2}$ and $e^{-r^{2}/w^{2}}$ with $w=35$ µm, respectively.
@@
Figure 8 (a) represent the complex amplitude distribution of the PAO. The ideal edge to compare with edge enhancement of the object by vortex phases with the amplitude function $1/r^{2}$ and $e^{-r^{2}/w^{2}}$ is same as Figs. 6 (c) and (d) but with a different pixel size i.e., $50\times 50$, which represents the total pixel size for the single-pixel edge enhancement reconstruction. The best edge enhancement using localized vortex phases with the amplitude function $1/r^{2}$ and $e^{-r^{2}/w^{2}}$ are shown in Figs. 8 (b) and (c), respectively. The amplitude function $e^{-r^{2}/w^{2}}$ could successfully enhance the edge of the object using the proposed single-pixel edge enhancement technique while the amplitude function $1/r^{2}$, which could successfully enhance the edge in Section 3.1, failed to enhance the edge. The main reason of this failure is considered as low fidelity (0.02), as shown in Table 1. Figure 9 shows dependence of the correlation coefficient between the single-pixel edge enhancement by using the Gaussian amplitude function and the ideal edge enhancement on the optimizing parameters $w$ and $c$. We selected the starting point of $w$ greater than 15 µm, which exceeds the radius of the super-pixel size of 10 µm, because $w$ values lower than 15 µm do not detect the edges of the object. The highest correlation coefficient 0.95 is achieved when $w=35$ µm and $c=6$. The edge enhancement using these optimized parameters is illustrated in Fig. 8 (c). The strong and positive relationship between the ideal edge and the proposed method’s edge demonstrates that the proposed approach is indeed capable of effectively enhancing the edges of objects.
@@ 4 Conclusion
In this paper, we proposed and numerically demonstrated edge enhancement of objects using single-pixel imaging via convolutional filtering, in which the localized vortex phase consists of a spiral phase $e^{{\mathrm{i}}\theta}$ and a localized amplitude function $a(\bm{r})$. We have tested several localized functions to find the highest fidelity using super-pixel method for use in our proposed method and a Gaussian amplitude function offers the highest fidelity (0.89). We numerically demonstrated the proposed method and the correlation coefficient between ideal edge enhancement and our method was 0.95. The results show that the proposed method is capable of enhancing the edge of the PAO while eliminating extra single-pixel measurements required by previous studies. Our proposed method would be applied to invisible wavelength range, such as near-infrared fluorescence and electronic circuit inspection through silicon semiconductor. We plan to conduct experimental demonstration of the proposed method in the future.
[…]
Funding Japan Society for the Promotion of Science (18KK0079, 20K05364); Research Foundation for Opto-Science and Technology.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.

### v2

@@ 3.2 Fidelity estimation of super-pixel method to generate localized vortex phase
The modulation accuracy of the super-pixel method between target fields and desired fields can be quantified by calculating the fidelity, $F$, as shown below[29]:
$F=\frac{|\bm{D}_{\text{desired}}^{*}\cdot\bm{D}_{\text{out}}|^{2}}{|\bm{D}_{\text{desired}}|^{2}|\bm{D}_{\text{out}}|^{2}},$ (5)
where * denotes the complex conjugate. The $\bm{D}_{\text{out}}$ and $\bm{D}_{\text{desired}}$ are one column vectors reshaped from the two-dimensional target fields, $E_{\text{out}}(\bm{r})$, and desired fields, $E_{\text{desired}}(\bm{r})$, respectively. When $F=1$, it indicates that the similarity between target fields and desired fields is the same, whereas when $F=0$, it means they are completely different.
We conduct a numerical simulation to estimate the fidelity between the target field generated by the super-pixel method and the desired localized vortex phase with the amplitude functions, $(c/r)^{2}$, $c/r$, $\sqrt{c/r}$ and $ce^{-r^{2}/w^{2}}$, where $c$ denotes the amplitude coefficient. The optimizing parameters to obtain the best fidelity in the super-pixel method are the amplitude coefficient $c$ and the aperture radius $R$. The optimal values were determined by a systematic parameter sweep: the fidelity $F$ was calculated for a wide range of each parameter $R/\rho$ from 0 to 4 and for the Gaussian function, the waist $w$ was also varied from 0 to 100 µm. The combination of values yielding the maximum fidelity $F$ was selected. We use $w=100$ µm, which gives the highest fidelity of $0.89$. The simulation conditions remain the same as those in Section 3.1, with an additional parameter, a focal length of the lens setting at 300 mm.
Figures 7 (a-d) show the simulation results of the fidelity between the ideal and the generated localized vortex phases with different $a(\bm{r})$: $(c/r)^{2}$, $c/r$, $\sqrt{c/r}$ and $ce^{-r^{2}/w^{2}}$, respectively. The horizontal and vertical axes in Fig. 7 show the aperture radius $R$ normalized by aperture position $\rho$ and the amplitude coefficient $c$ of $h(\bm{r})$, respectively. Table 1 shows a summary of the maximum fidelities and the corresponding optimized parameters. The Gaussian amplitude function $e^{-r^{2}/w^{2}}$ provides a fidelity of $0.89$, whereas $\frac{1}{r^{2}}$, $\frac{1}{r}$ and $\frac{1}{\sqrt{r}}$ yield fidelity values below $0.6$ (refer to Table 1), because these power functions increase rapidly as they approach the origin, resulting in low fidelity when implemented with the finite amplitude range of the super-pixel method. Even if the fidelity to the desired function is low, however, edge enhancement might be possible if a certain degree of localization is achieved by the amplitude function. Thus, in the next section, we use the amplitude functions $\frac{1}{r^{2}}$ and $e^{-r^{2}/w^{2}}$ which successfully perform edge enhancement in Fig. 6, to evaluate our single-pixel method via convolutional filtering.
[…]
@@
Parameter $a(\bm{r})$ $\frac{1}{r^{2}}$ $\frac{1}{r}$ $\frac{1}{\sqrt{r}}$ $e^{-r^{2}/w^{2}}$
Max. fidelity, $F$ 0.02 0.45 0.57 0.89
Radius, $R/\rho$ 2.5 2.6 2 0.7
Amp. coefficient, $c$ 42 µm 100 µm 100 µm 9
@@
Now we evaluate the edge enhancement capability of our proposed single-pixel imaging method with the localized amplitude functions, $\frac{1}{r^{2}}$ and $e^{-r^{2}/w^{2}}$. We first prepared the binary hologram on the DMD to generate the shifted localized vortex phase $h(\bm{r}^{\prime}-\bm{r})$ by using the super-pixel method with $n=4$. By following the principles of Section 2.1, we calculated the Fourier transform of the object light multiplied with the hologram on the DMD and then obtained optical power at the center point of the aperture, $(x,y)=(-\rho,n\rho)$, to reconstruct the two-dimensional edge enhanced image. The simulation conditions remained the same as those in Section 3.1.
Figure 8: (a) Distribution of PAO, (b) and (c) Edge enhancement via proposed method using the optical vortex with the amplitude function $1/r^{2}$ and $e^{-r^{2}/w^{2}}$ with $w=35$ µm, respectively.
@@
Figure 8 (a) represents the complex amplitude distribution of the PAO. The ideal edge to compare with the edge enhancement of the object by vortex phases with the amplitude function $1/r^{2}$ and $e^{-r^{2}/w^{2}}$ is the same as Fig. 6 (b) but with a different pixel size i.e., $50\times 50$, which represents the total pixel size for the single-pixel edge enhancement reconstruction. The best edge enhancement using localized vortex phases with the amplitude function $1/r^{2}$ and $e^{-r^{2}/w^{2}}$ are shown in Figs. 8 (b) and (c), respectively. The amplitude function $e^{-r^{2}/w^{2}}$ could successfully enhance the edge of the object using the proposed single-pixel edge enhancement technique, while the amplitude function $1/r^{2}$, which could successfully enhance the edge in Section 3.1, failed to enhance the edge. The main reason of this failure is considered to be low fidelity (0.02), as shown in Table 1. Figure 9 shows the dependence of the correlation coefficient between the single-pixel edge enhancement by using the Gaussian amplitude function and the ideal edge enhancement on the optimizing parameters $w$ and $c$. We selected a starting point of $w$ greater than 15 µm, which exceeded the radius of the super-pixel size of 10 µm, because $w$ values lower than 15 µm did not detect the edges of the object. The highest correlation coefficient 0.95 was achieved when $w=35$ µm and $c=6$. The edge enhancement using these optimized parameters is illustrated in Fig. 8 (c). The strong and positive relationship between the ideal edge and the edge produced by the proposed method demonstrates its effectiveness in enhancing the edges of objects.
@@ 4 Conclusion
In this paper, we proposed and numerically demonstrated the edge enhancement of objects using single-pixel imaging via convolutional filtering with a localized vortex phase (spiral phase $e^{i\theta}$ and amplitude function $a(\bm{r})$). Through systematic testing of localized functions, the Gaussian amplitude achieved the highest fidelity (0.89) using the super-pixel method. Our simulations yielded a 0.95 correlation coefficient between ideal and enhanced edges, confirming the method’s ability to enhance phase-amplitude object (PAO) edges while eliminating the measurement overhead of phase-shifting approaches. This method shows strong potential for non-visible wavelength applications such as near-infrared fluorescence microscopy and silicon semiconductor circuit inspection, where conventional sensors face limitations.
Future work will focus on three key directions to further advance this methodology: (1) validation across diverse samples including biological specimens and semiconductor inspection patterns to establish robustness in real-world applications, (2) optimization of vortex generation through comparative analysis of the super-pixel method with alternative approaches such as error diffusion [27] and Direct Binary Search [28] to evaluate speed-fidelity trade-offs, and (3) single-pixel edge enhancement with randomly-distributed localized vortex filter for faster acquisition time.
Declarations
  • Funding
    Japan Society for the Promotion of Science (18KK0079, 20K05364); Research Foundation for Opto-Science and Technology.
  • Conflict of interest/Competing interests
    The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
  • Ethics approval and consent to participate
    Not applicable.
  • Consent for publication
    Not applicable.
  • Data availability
    Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
  • Author contribution
    All authors contributed equally.